OpenStudy (anonymous):

Is there a simple way to find the zeros of x^6-1?

geerky42 (geerky42):

Let x³ = u \(\Large u^2 = (x^3)^2 = x^6\) u²-1 = 0 (u+1)(u-1)=0 u = ±1 x³ = ±1 x = ³√(±1) = ±1 = \(\boxed{\text{1 or -1}}\)

geerky42 (geerky42):

That's my approach. Is this clear?

OpenStudy (anonymous):

That's a perfect solution @geerky42 @sfb Do you understand what geerky has done here?

OpenStudy (anonymous):

No, I don't think so. I'm still trying to understand it.

geerky42 (geerky42):

What part don't you understand? I'd be glad to clarify it.

OpenStudy (anonymous):

Okay, I got x=+/-1, x=0, x=+/-i

OpenStudy (anonymous):

Would that be right?

geerky42 (geerky42):

x is equal only to ±1. Because if you plug in 0 and ±i into x⁶-1=0, it won't be equal to zero.

OpenStudy (anonymous):

I originally got +/-1, but my teacher said I needed to find the 4 complex zeros.

geerky42 (geerky42):

@genius12 Can you help her? I'm not good with complex...

OpenStudy (anonymous):

4 complex zeros for x^6 - 1? It only has 2 zeros both of which are complex. When I say they are complex I mean they are real numbers, but because real numbers are a subset of complex numbers, then any real roots are complex roots. What makes your teacher think that they're 4 solutions?

OpenStudy (anonymous):

"These are the correct real zeros, but you need to find the 4 complex zeros. Hint: factor (x6 - 1) into (x3 + 1)(x3 - 1), then keep factoring or solving to find the other complex zeros."

OpenStudy (anonymous):

Oh lol. I see what you mean now.

geerky42 (geerky42):

I think I saw my mistakes, ³√(±1) is COMPLEX, not real. so far we have found two complex zeros.

geerky42 (geerky42):

According to Wolfram Alpha, x⁶-1 = (x-1) (x+1) (x²-x+1) (x²+x+1).

OpenStudy (anonymous):

I would use the quadratic formula on (x^2-x+1) and (x²+x+1)?

OpenStudy (anonymous):

That's how I got my i's.

geerky42 (geerky42):

"I would use the quadratic formula on (x^2-x+1) and (x²+x+1)?" Yes.

OpenStudy (anonymous):

Shoot. I screwed. Ya, do the factorisation like Wolframalpha shows and then use quadratic formula to factor the quadratics to get the complex roots.

OpenStudy (anonymous):

Haha, okay. Thank you both for the help! It is much appreciated.

geerky42 (geerky42):

You're welcome.